determine the fractional change in volume as the pressure of the atmosphere 1.0×10^5 pa around a metal block is reduced to zero by placing the block in vacuum the bulk modulus for the block is 1.25×10^11
1
Expert's answer
2018-08-29T12:55:08-0400
Bulk Modulus
K=(-∆P)/((∆V/V) )
where ∆P is change in pressure and ∆ V is change in volume.
So,
∆V/V=(-∆P)/K
Substituting,
∆V/V=(-(0-1.0×〖10〗^5))/(1.25×〖10〗^11 )=8×〖10〗^(-7)